OpenStudy (anonymous):

Find the equation in standard form which passes through (–2, –7) and (1, –1).

hero (hero):

Find the slope, then use point slope form

OpenStudy (anonymous):

y = mx +b ?

OpenStudy (anonymous):

The slope is 2.

hero (hero):

Yes, correct

hero (hero):

m(x - x_1) = (y - y_1) 2(x - 1) = (y + 1) 2x - 2 = y + 1 y = 2x - 3

OpenStudy (anonymous):

I just got that (: lol Thank you

OpenStudy (jamesj):

No, the slope is NOT 2.

hero (hero):

\[m = \frac{y_2 - y_1}{x_2 - x_1} = \frac{-7 - (-1)}{-2 - 1} = \frac{-7 + 1}{-3} = \frac{-6}{-3}=2 \]

hero (hero):

The slope IS REALLY 2